# Towards a generic absoluteness theorem for Chang models

### v2

@@ 4. The UB-Capturing Principle
Our initial idea, which originated in [22], was to use an $\omega$-sequence of countable substructures of $V_{\xi}$, which we were calling blocks. Our goal was to iterate these $\omega$ many models simultaneously to make the reals generic in such a way that $L(\Gamma^{\infty},\mathbb{R})$ could be realized as the derived model of the final iterate. This approach appeared in [22]. This approach presents the models at the beginning of the process, each model on the sequence corresponds to the “changing the model” step applied to the previous model on the sequence.
@@
    Suppose further that $\gamma$ is an inaccessible cardinal. Then if $\xi\in X\cap\gamma$ and $Y\subseteq V_{\xi}$ then $\sup(X\cap\gamma)=\sup(X(Y)\cap\gamma)$ and $(X\cap V_{\gamma})(Y)\prec V_{\gamma}$.
@@
To show the first claim, it is enough to see that $X[a]\cap V_{\gamma}\subseteq Y$ given that $a\in Y$. Let $\xi$ be the rank of $a$ and let $\zeta\in X\cap\gamma$ be bigger than $\xi$, which exists because $\sup(X\cap\gamma)=\sup(Y\cap\gamma)$ and $\xi\in Y$. Suppose now that $f\in X$, $f(a)\in V_{\gamma}$ and we want to see that $f(a)\in Y$. We then have that $f\restriction V_{\zeta}\in X\cap V_{\gamma}\subseteq Y$ and hence, $f(a)=(f\restriction V_{\zeta})(a)\in Y$.
To show the second clause, it is enough to show the claim holds for $Y=\{a\}$. Suppose then that $\gamma$ is an inaccessible cardinal, $\xi\in X\cap\gamma$ and $a\in V_{\xi}$. We want to see that if $f\in X$ is such that $f(a)$ is an ordinal then $f(a)<\sup(X\cap\gamma)$. Let $A\subseteq V_{\xi}$ be the set of those $u\in V_{\xi}$ such that $f(u)$ is an ordinal. Then $A\in X$. Let then $\tau=\sup\{f(u):u\in A\}$. Because $\gamma$ is inaccessible we have that $\tau<\gamma$. Since clearly $\tau\in X$, we have that $f(a)<\sup(X\cap\gamma)$.
Finally, suppose $\gamma$ is an inaccessible cardinal, $\xi\in X\cap\gamma$ and $a\in V_{\xi}$. We want to see that $(X\cap V_{\gamma})[a]\prec V_{\gamma}$. This is because $X[a]\cap V_{\gamma}=(X\cap V_{\gamma})[a]$. Indeed, let $f\in X$ be such that $f(a)\in V_{\gamma}$ and let $A\subseteq V_{\xi}$ be the set of those $u$ such that $f(u)\in V_{\gamma}$. Then $A\in X$ and hence, $A\in X\cap V_{\gamma}$. Therefore, $f\restriction A\in X\cap V_{\gamma}$. But we have that $f(a)=f\restriction A(a)$, implying that $f(a)\in(X\cap V_{\gamma})[a]$. ∎
@@
Suppose $\chi>\eta$ are inaccessible cardinals, $W\in V_{\chi}$, $X\prec V_{\chi}$ is a strong substructure of $V_{\chi}$ and $Y\prec V_{\chi}$. We say that $X$ is $(W\mid\eta)$-full if $\eta\in X$ and for every $b\in V_{\eta}$,
@@
The above concepts also make sense in generic extensions of $V$.
@@
Suppose $\chi>\eta$ are inaccessible cardinals, $g$ is $<\eta$-generic, $A\in\Gamma^{\infty}_{g}$, $X\prec V_{\chi}$ with $X\in V[g]$ and $W\in V_{\chi}$ with ${\sf{comp}}(W)>\eta$. We say $(X,W)$ captures $A$ if there is a $\bar{\mu}\subseteq X\cap W$ such that (in $V[g]$) $A=S_{\bar{\mu}_{g}}$.
We are now ready to introduce the UB-Capturing Principle. If $W$ is any set then we let ${\sf{comp}}(W)$ be the largest cardinal $\kappa$ such that every ultrafilter in $W$ is $\kappa$-complete.
@@
Suppose the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(W,X)$. Let $h$ be ${<}\lambda$-generic, $A\in\Gamma^{\infty}_{h}$, $z\subseteq W$ witness Clause 4 of Definition 4.7 applied to $(h,A)$ and let $y\in V_{\eta}[h]\cap{\wp}(V_{\eta})$. Set $Y=X(y)$ and $Z=Y(z)$, and let $M_{Y}$ be the transitive collapse of $Y$ and $M_{Z}$ be the transitive collapse of $Z$. Let $\pi_{Y}:M_{Y}\rightarrow V_{\gamma}$ and $\pi_{Z}:M_{Z}\rightarrow V_{\gamma}$ be the inverses of the transitive collapses, and let $\pi_{Y,Z}=(\pi_{Z})^{-1}\circ\pi_{Y}$. Then $\pi_{Y}$ and $\pi_{Z}$ are elementary, ${\rm cp}(\pi_{Y,Z})>\pi_{Y}^{-1}(\eta)$ and $Z$ captures $A$.
@@
Clause 2 of Lemma 4.3 and the fact that $W\in X$ imply that $Y$ and $Z$ are elementary substructures of $V_{\chi}$, and hence, $\pi_{Y}$ and $\pi_{Z}$ are elementary. Clearly, since $X(z)$ captures $A$, $Z=Y(z)$ captures $A$.
@@
It is not hard to show that provided there is a class of Woodin cardinals, if $\eta\geq\lambda$ are inaccessible cardinals then there is a $(\lambda,\eta)$-saturated $W$. This follows from Theorem 2.3. We can now prove our main theorem on UB-Capturing Principle.
@@
First, since $X\prec V_{\chi+1}$, we have that $j[X\cap V_{\chi}]\prec M_{j(\chi)}$ and $j[X\cap V_{\chi}]$ is a strong substructure of $M_{j(\chi)}$. Next since $j\restriction X:X\rightarrow j[X]$ is a surjection and ${|X|}^{V}<\lambda$, ${|(j[X])|}^{M}<\lambda$.
We now list some facts which we will need to establish Clause 3 of Definition 4.7. We clearly have that
@@
(2) then implies Clause 4 of Definition 4.7. Clause 3 is witnessed by $j(W)$. It remains to prove Clause 2 of Definition 4.7. Notice first that we have that $j(\kappa)\in j[X\cap V_{\chi}]$. We thus need to show that in $M$, $j[X\cap V_{\chi}]$ is $(j[W]\mid j(\kappa))$-full. To see this, it is enough to prove the following claim.
@@
In what follows we will describe how to realize $L(\Gamma^{\infty},\mathbb{R})$ as the derived model of the direct limit model of a linear system of models. Each model in this system will be obtained via the one step construction described in the next section. The construction of the entire linear system will be performed in $V[g]$ where $g$ is a generic enumeration of $\Gamma^{\infty}$ in order type $\omega$ or $\omega_{1}$. Let $\iota\in\{\omega,\omega_{1}\}$ and let $(M_{\alpha},n_{\alpha,\beta}\mid\alpha<\beta<\iota)$ be the linear sequence our construction produces. Also, let $(A_{\alpha}\mid\alpha<\iota)$ be the generic enumeration of $\Gamma^{\infty}$. We will have an inaccessible cardinal $\chi$ and embeddings $\pi_{\alpha}:M_{\alpha}\rightarrow V_{\chi}$ such that for $\alpha<\beta<\iota$, $\pi_{\alpha}=\pi_{\beta}\circ n_{\alpha,\beta}$. The one step construction is a construction that describes the procedure for obtaining $M_{\alpha+1}$ from $M_{\alpha}$. Typically, this will have two parts. The first part produces an intermediate model $N$ from $M_{\alpha}$ along with an embedding $\sigma:N\rightarrow V_{\chi}$ such that $\operatorname{rng}(\pi_{\alpha})\subseteq\operatorname{rng}(\sigma)$. $N$’s job is to ensure that $A_{\alpha}$ will be in the derived model of $M_{\iota}$, the direct limit of $(M_{\alpha},n_{\alpha,\beta}\mid\alpha<\beta<\iota)$. The construction of $N$ will involve genericity iterations that make homogeneously Suslin representations of $A_{\alpha}$ generic. The second part will be executed by appealing to Lemma 4.8. This part will produce $M_{\alpha+1}$ from $N$ using Lemma 4.8, which we will apply to $Y=_{def}\sigma[N]$ and $A_{\alpha+1}$. The $M_{Z}$ of that lemma will be our $M_{\alpha+1}$. The entire sequence will be produced by consecutively applying the one step construction and taking direct limits at limit stages. In the next section, we will describe the one step construction.
@@ 5.1. Martin-Steel Iterability Theorem
If ${\mathcal{T}}$ is an iteration tree then ${\sf gen}({\mathcal{T}})$ is the supremum of lengths of the extenders used in ${\mathcal{T}}$. First we isolate the models that we will iterate.
@@
    $M^{\sf d}\models``E$ is a nice extender” [^20] I.e., $\operatorname{lh}(E)$ is also the strength of $E$ and $\operatorname{lh}(E)$ is an inaccessible cardinal.
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock. Suppose that ${\mathcal{T}}$ is a normal iteration tree on ${\sf a}$ such that $\operatorname{lh}({\mathcal{T}})=\omega$. There is then a well-founded branch $b$ of ${\mathcal{T}}$ and an embedding $k:M^{\mathcal{T}}_{b}\rightarrow U_{\chi}$ such that $(k,M^{\mathcal{T}}_{b},\kappa^{\sf a})$ is an iterate of ${\sf a}$ via ${\mathcal{T}}^{\frown}\{b\}$ (so that $\sigma^{\sf a}=k\circ\pi^{\mathcal{T}}_{b}$).
@@
Let $\pi:W\rightarrow V_{\chi+1}$ be an elementary embedding with $W$ countable (in $V$) and transitive and such that $U_{\chi}\in\operatorname{rng}(\pi)$ and $\operatorname{rng}(\sigma^{\sf a})\subseteq\operatorname{rng}(\pi)$. Set $N=(\pi)^{-1}(U_{\chi})$ and $m=\pi\restriction N$. Then there is $n:M^{{\sf a}}\rightarrow N$ such that $\sigma^{\sf a}=m\circ n$. Let now ${\mathcal{U}}$ be the $n$-copy of ${\mathcal{T}}$. Notice that because ${\mathcal{U}}$ is above $m^{-1}(\lambda)$ [^22] We have that ${\mathcal{T}}$ is above $\kappa^{\sf a}$ and $\sigma^{\sf a}(\kappa^{\sf a})>\lambda$. Thus, ${\mathcal{U}}$ is above $m^{-1}(\lambda)$., ${\mathcal{U}}$ lifts to an iteration tree ${\mathcal{U}}^{\prime}$ on $W$. This is because $(\pi)^{-1}(g)$ is a small forcing relative to the critical points of extenders used in ${\mathcal{U}}$ [^23] Recall that our iteration trees are nice, see Definition 5.1. These are the trees used in [14]. See [14, Remark 2.1].. It now follows from Martin-Steel Iterability Theorem ([14, Theorem 2.3]) that there is a branch $b$ of ${\mathcal{U}}^{\prime}$ and an elementary embedding $k^{\prime}:M^{{\mathcal{U}}^{\prime}}_{b}\rightarrow V_{\chi+1}$ such that $\pi=k^{\prime}\circ\pi^{{\mathcal{U}}^{\prime}}_{b}$.
A careful examination of the relationship between ${\mathcal{U}}$ and ${\mathcal{U}}^{\prime}$ shows that if $h=(\pi)^{-1}(g)$ then $\pi^{{\mathcal{U}}^{\prime}}_{b}\restriction N[h]$ is simply the lift-up of $\pi^{\mathcal{U}}_{b}$ to $N[h]$. More precisely, $\pi^{{\mathcal{U}}^{\prime}}_{b}(\tau_{h})=(\pi^{\mathcal{U}}_{b}(\tau))_{h}$ where $\mu_{h}$ is the realization of the name $\mu$ by $h$.
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock. Suppose $E\in M^{\sf a}$ is a nice extender such that $\sigma^{\sf a}({\rm cp}(E))\geq\lambda$. There is then an elementary embedding $k:Ult(M^{\sf a},E)\rightarrow U_{\chi}$ such that $\sigma^{\sf a}=k\circ\pi_{E}^{M^{\sf a}}$ [^24] Equivalently, ${\sf{b}}=(k,Ult(M^{\sf a},E),\pi_{E}(\kappa^{\sf a}))$ is an ${\sf a}$-realizable rock such that $\pi_{E}^{M^{\sf a}}=\sigma^{{\sf a},{\sf{b}}}$..
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and $\pi:M\rightarrow U_{\chi}$ is an elementary embedding with $M$ transitive and countable in $V$. Suppose $E$ is an $M$-extender such that $\pi({\rm cp}(E))>\lambda$. We say that $E$ is derivable from $\pi$ if there is $\xi<\chi$ and an injective order-preserving map $m:\operatorname{lh}(E)\rightarrow\xi$ such that for every $a\in\operatorname{lh}(E)^{<\omega}$ and for every $A\in{\wp}([{\rm cp}(E)]^{{|a|}})\cap M$, $(a,A)\in E$ if and only if $m[a]\in\pi(A)$.
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$), ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock and ${\sf{b}}$ is an ${\sf a}$-realizable rock such that ${\rm cp}(\sigma^{{\sf a},{\sf{b}}})>\kappa^{{\sf a}}$ (thus, $\kappa^{\sf a}=\kappa^{\sf{b}}$). Suppose $E\in M^{\sf{b}}$ is a nice extender with ${\rm cp}(E)=\kappa^{\sf a}$. Let $k:Ult(M^{\sf{b}},E)\rightarrow U_{\chi}$ be as in Lemma 5.8 (so $\sigma^{{\sf{b}}}=k\circ\pi_{E}^{M^{\sf{b}}}$) and set $\xi=k(\operatorname{lh}(E))$ and $m=k\restriction\operatorname{lh}(E)$. Then

### v3

@@ 4. The UB-Capturing Principle
Our initial idea, which originated in [21], was to use an $\omega$-sequence of countable substructures of $V_{\xi}$, which we were calling blocks. Our goal was to iterate these $\omega$ many models simultaneously to make the reals generic in such a way that $L(\Gamma^{\infty},\mathbb{R})$ could be realized as the derived model of the final iterate. This approach appeared in [21]. It presents the models at the beginning of the process, each model on the sequence corresponds to the “changing the model” step applied to the previous model on the sequence.
@@
    Suppose further that $\gamma$ is an inaccessible cardinal. If $\xi\in X\cap\gamma$ and $Y\subseteq V_{\xi}$, then $\sup(X\cap\gamma)=\sup(X(Y)\cap\gamma)$ and $(X\cap V_{\gamma})(Y)\prec V_{\gamma}$.
@@
To show the first claim, it suffices to see that $X[a]\cap V_{\gamma}\subseteq Y$ given that $a\in Y$. Let $\xi$ be the rank of $a$ and let $\zeta\in X\cap\gamma$ be larger than $\xi$, which exists because $\sup(X\cap\gamma)=\sup(Y\cap\gamma)$ and $\xi\in Y$. Suppose now that $f\in X$, $f(a)\in V_{\gamma}$ and we want to see that $f(a)\in Y$. We then have that $f\restriction V_{\zeta}\in X\cap V_{\gamma}\subseteq Y$ and hence, $f(a)=(f\restriction V_{\zeta})(a)\in Y$.
To show the second clause, it is enough to show the claim holds for $Y=\{a\}$. Suppose then that $\gamma$ is an inaccessible cardinal, $\xi\in X\cap\gamma$ and $a\in V_{\xi}$. We want to see that if $f\in X$ is such that $f(a)$ is ordinal, then $f(a)<\sup(X\cap\gamma)$. Let $A\subseteq V_{\xi}$ be the set of those $u\in V_{\xi}$ such that $f(u)$ is an ordinal. Then $A\in X$. Let then $\tau=\sup\{f(u):u\in A\}$. Because $\gamma$ is inaccessible we have $\tau<\gamma$. Since clearly $\tau\in X$, we have $f(a)<\sup(X\cap\gamma)$.
Finally, suppose $\gamma$ is an inaccessible cardinal, $\xi\in X\cap\gamma$ and $a\in V_{\xi}$. We want to see that $(X\cap V_{\gamma})[a]\prec V_{\gamma}$. This is because $X[a]\cap V_{\gamma}=(X\cap V_{\gamma})[a]$. Indeed, let $f\in X$ be such that $f(a)\in V_{\gamma}$ and let $A\subseteq V_{\xi}$ be the set of those $u$ such that $f(u)\in V_{\gamma}$. Then $A\in X$ and hence, $A\in X\cap V_{\gamma}$. Therefore, $f\restriction A\in X\cap V_{\gamma}$. But we have $f(a)=f\restriction A(a)$, which implies that $f(a)\in(X\cap V_{\gamma})[a]$. ∎
@@
Suppose $\chi>\eta$ are inaccessible cardinals, $W\in V_{\chi}$, and $X\prec V_{\chi}$ is a strong substructure of $V_{\chi}$. We say that $X$ is $(W\mid\eta)$-full if $\eta\in X$ and for every $b\in V_{\eta}$,
@@
The above concepts also make sense in generic extensions of $V$. If $W$ is any set then we let ${\sf{comp}}(W)$ be the largest cardinal $\kappa$ such that every ultrafilter in $W$ is $\kappa$-complete.
@@
Suppose $\chi>\eta$ are inaccessible cardinals, $g$ is ${<}\eta$-generic, $A\in\Gamma^{\infty}_{g}$, $X\prec V_{\chi}$ with $X\in V[g]$ and $W\in V_{\chi}$ with ${\sf{comp}}(W)>\eta$. We say $(X,W)$ captures $A$ if there is a homogeneity system $\bar{\mu}\subseteq X\cap W$ such that (in $V[g]$) $A=S_{\bar{\mu}_{g}}$.
If there is a $W$ such that $(X,W)$ captures $A$ we also just say that $X$ captures $A$. We are now ready to introduce the UB-Capturing Principle.
@@
Suppose the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(W,X)$. Let $h$ be ${<}\lambda$-generic, $A\in\Gamma^{\infty}_{h}$, $z\subseteq W$ witness Clause (4) of Definition 4.7 applied to $(h,A)$ and let $y\in V_{\eta}[h]\cap{\wp}(V_{\eta})$. Set $Y=X(y)$ and $Z=Y(z)$, and let $M_{Y}$ be the transitive collapse of $Y$ and $M_{Z}$ be the transitive collapse of $Z$. Let $\pi_{Y}:M_{Y}\rightarrow V_{\gamma}$ and $\pi_{Z}:M_{Z}\rightarrow V_{\gamma}$ be the inverses of the transitive collapses, and let $\pi_{Y,Z}=(\pi_{Z})^{-1}\circ\pi_{Y}$. Then $\pi_{Y}$ and $\pi_{Z}$ are elementary, ${\rm cp}(\pi_{Y,Z})>\pi_{Y}^{-1}(\eta)$ and $Z$ captures $A$.
@@
Clause (2) of Lemma 4.3 and the fact that $W\in X$ imply that $Y$ and $Z$ are elementary substructures of $V_{\chi}$, and hence, $\pi_{Y}$ and $\pi_{Z}$ are elementary. Clearly, since $X(z)$ captures $A$, $Z=Y(z)$ captures $A$.
@@
It is not hard to show that provided there is a class of Woodin cardinals, if $\eta\geq\lambda$ are inaccessible cardinals then there is a $(\lambda,\eta)$-saturated $W$. This follows from Theorem 2.3. We can now prove our main theorem on the UB-Capturing Principle.
@@
First, since $X\prec V_{\chi+1}$, we have that $j[X\cap V_{\chi}]\prec M_{j(\chi)}$ and $j[X\cap V_{\chi}]$ is a strong substructure of $M_{j(\chi)}$ (see Clause (2) of Lemma 4.3). Next, since $j\restriction X:X\rightarrow j[X]$ is a surjection and ${|X|}^{V}<\lambda$, ${|(j[X])|}^{M}<\lambda$.
We now list some facts which we will need to establish Clause (3) of Definition 4.7. We clearly have that
@@
(2) then implies Clause (4) of Definition 4.7. Clause (3) of Definition 4.7 is witnessed by $j(W)$. It remains to prove Clause (2) of Definition 4.7. Notice first that we have that $j(\kappa)\in j[X\cap V_{\chi}]$. We thus need to show that in $M$, $j[X\cap V_{\chi}]$ is $(j[W]\mid j(\kappa))$-full. To see this, it is enough to prove the following claim.
@@
This finishes the proof of Theorem 4.10. ∎
In what follows we will describe how to realize $L(\Gamma^{\infty},\mathbb{R})$ as the derived model of the direct limit model of a linear system of models. Each model in this system will be obtained via the one step construction described in the next section. The construction of the entire linear system will be performed in $V[g]$ where $g$ is a generic enumeration of $\Gamma^{\infty}$ in order type $\omega$ or $\omega_{1}$. Let $\iota\in\{\omega,\omega_{1}\}$ and let $(M_{\alpha},n_{\alpha,\beta}\mid\alpha<\beta<\iota)$ be the linear sequence our construction produces with $n_{\alpha,\beta}:M_{\alpha}\rightarrow M_{\beta}$. Also, let $(A_{\alpha}\mid\alpha<\iota)$ be the generic enumeration of $\Gamma^{\infty}$. We will have an inaccessible cardinal $\chi$ and embeddings $\pi_{\alpha}:M_{\alpha}\rightarrow V_{\chi}$ such that for $\alpha<\beta<\iota$, $\pi_{\alpha}=\pi_{\beta}\circ n_{\alpha,\beta}$. The one step construction is a construction that describes the procedure for obtaining $M_{\alpha+1}$ from $M_{\alpha}$. Typically, this will have two parts. The first part produces an intermediate model $N$ from $M_{\alpha}$ along with an embedding $\sigma:N\rightarrow V_{\chi}$ such that $\operatorname{rng}(\pi_{\alpha})\subseteq\operatorname{rng}(\sigma)$. $N$’s job is to ensure that $A_{\alpha}$ will be in the derived model of $M_{\iota}$, the direct limit of $(M_{\alpha},n_{\alpha,\beta}\mid\alpha<\beta<\iota)$. The construction of $N$ will involve genericity iterations that make homogeneously Suslin representations of $A_{\alpha}$ generic. The second part will be executed by appealing to Lemma 4.8. This part will produce $M_{\alpha+1}$ from $N$ using Lemma 4.8, which we will apply to $Y=_{def}\sigma[N]$ and $A_{\alpha+1}$. The $M_{Z}$ of that lemma will be our $M_{\alpha+1}$. The entire sequence will be produced by consecutively applying the one step construction and taking direct limits at limit stages. In the next section, we will describe the one step construction in more detail.
@@ 5.1. Martin-Steel Iterability Theorem
If ${\mathcal{T}}$ is an iteration tree then ${\sf gen}({\mathcal{T}})$ is the supremum of the lengths of the extenders used in ${\mathcal{T}}$. First we isolate the models that we will iterate.
@@
    $M^{\sf d}\models``E$ is a nice extender”
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock. Suppose that ${\mathcal{T}}$ is a normal iteration tree on ${\sf a}$ such that $\operatorname{lh}({\mathcal{T}})=\omega$ [^20] This means that ${\mathcal{T}}$ is above $\kappa^{\sf a}$. There is then a well-founded branch $b$ of ${\mathcal{T}}$ and an embedding $k:M^{\mathcal{T}}_{b}\rightarrow U_{\chi}$ such that $(k,M^{\mathcal{T}}_{b},\kappa^{\sf a})$ is an iterate of ${\sf a}$ via ${\mathcal{T}}^{\frown}\{b\}$ (so that $\sigma^{\sf a}=k\circ\pi^{\mathcal{T}}_{b}$).
@@
Let $\pi:W\rightarrow V_{\chi+1}$ be an elementary embedding with $W$ countable (in $V$) and transitive and such that $U_{\chi}\in\operatorname{rng}(\pi)$ and $\operatorname{rng}(\sigma^{\sf a})\subseteq\operatorname{rng}(\pi)$. Set $N=(\pi)^{-1}(U_{\chi})$ and $m=\pi\restriction N$. Then there is $n:M^{{\sf a}}\rightarrow N$ such that $\sigma^{\sf a}=m\circ n$. Let now ${\mathcal{U}}$ be the $n$-copy of ${\mathcal{T}}$. Notice that because ${\mathcal{U}}$ is above $m^{-1}(\lambda)$ [^21] We have that ${\mathcal{T}}$ is above $\kappa^{\sf a}$ and $\sigma^{\sf a}(\kappa^{\sf a})>\lambda$. Thus, ${\mathcal{U}}$ is above $m^{-1}(\lambda)$., ${\mathcal{U}}$ lifts to an iteration tree ${\mathcal{U}}^{\prime}$ on $W$. This is because $(\pi)^{-1}(g)$ is a small forcing relative to the critical points of extenders used in ${\mathcal{U}}$ [^22] Recall that our iteration trees are nice, see Definition 5.1. These are the trees used in [14]. See [14, Remark 2.1].. It now follows from the Martin-Steel Iterability Theorem ([14, Theorem 2.3]) that there is a branch $b$ of ${\mathcal{U}}^{\prime}$ and an elementary embedding $k^{\prime}:M^{{\mathcal{U}}^{\prime}}_{b}\rightarrow V_{\chi+1}$ such that $\pi=k^{\prime}\circ\pi^{{\mathcal{U}}^{\prime}}_{b}$.
A careful examination of the relationship between ${\mathcal{U}}$ and ${\mathcal{U}}^{\prime}$ shows that if $h=(\pi)^{-1}(g)$ then $\pi^{{\mathcal{U}}^{\prime}}_{b}\restriction N[h]$ is simply the lift-up of $\pi^{\mathcal{U}}_{b}$ to $N[h]$. More precisely, $\pi^{{\mathcal{U}}^{\prime}}_{b}(\tau_{h})=(\pi^{\mathcal{U}}_{b}(\tau))_{h}$ where $\tau_{h}$ and $(\pi^{\mathcal{U}}_{b}(\tau))_{h}$ are the realizations of the names $\tau$ and $\pi^{\mathcal{U}}_{b}(\tau)$ by $h$.
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and a $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock. Suppose $E\in M^{\sf a}$ is a nice extender such that $\sigma^{\sf a}({\rm cp}(E))\geq\lambda$. There is then an elementary embedding $k:Ult(M^{\sf a},E)\rightarrow U_{\chi}$ such that $\sigma^{\sf a}=k\circ\pi_{E}^{M^{\sf a}}$ [^23] Equivalently, ${\sf{b}}=(k,Ult(M^{\sf a},E),\pi_{E}(\kappa^{\sf a}))$ is an ${\sf a}$-realizable rock such that $\pi_{E}^{M^{\sf a}}=\sigma^{{\sf a},{\sf{b}}}$..
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and a $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$) and $\pi:M\rightarrow U_{\chi}$ is an elementary embedding with $M$ transitive and countable in $V$. Suppose $E$ is an $M$-extender such that $\pi({\rm cp}(E))>\lambda$. We say that $E$ is derivable from $\pi$ if there is $\xi<\chi$ and an injective order-preserving map $m:\operatorname{lh}(E)\rightarrow\xi$ such that for every $a\in\operatorname{lh}(E)^{<\omega}$ and for every $A\in{\wp}([{\rm cp}(E)]^{{|a|}})\cap M$, $(a,A)\in E$ if and only if $m[a]\in\pi(A)$.
@@
Suppose $\chi>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground of $V$ as witnessed by $\mathbb{P}\in U_{\lambda}$ and a $U$-generic $g\subseteq\mathbb{P}$ (thus, $V=U[g]$), ${\sf a}$ is a $(U,\chi,\lambda)$-realizable rock and ${\sf{b}}$ is an ${\sf a}$-realizable rock such that ${\rm cp}(\sigma^{{\sf a},{\sf{b}}})>\kappa^{{\sf a}}$ (thus, $\kappa^{\sf a}=\kappa^{\sf{b}}$). Suppose $E\in M^{\sf{b}}$ is a nice extender with ${\rm cp}(E)=\kappa^{\sf a}$. Let $k:Ult(M^{\sf{b}},E)\rightarrow U_{\chi}$ be as in Lemma 5.8 (so $\sigma^{{\sf{b}}}=k\circ\pi_{E}^{M^{\sf{b}}}$) and set $\xi=k(\operatorname{lh}(E))$ and $m=k\restriction\operatorname{lh}(E)$. Then